# Simulation of Polynomial-Time Nondeterministic Turing Machines via Universal Nondeterministic Turing Machines

### v18

@@ 1.3. Our Approach
As a novel idea and attempt, we will apply simulation techniques. To the best of our knowledge, if we want to create a language not in a specific complexity class $\mathcal{C}$, the most obvious way is to apply the diagonalization technique or the lazy-diagonalization technique, because the simulation and diagonalization technique [Tur37, HS65] is a standard method (currently the only way) to prove lower bounds on uniform computing models (i.e., the Turing machine model). But, we find that it is hard to apply standard diagonalization or lazy-diagonalization techniques to produce a language that not in ${\rm co}\mathcal{NP}$. Instead, by the relation between the language $L(M)$ accepted by polynomial-time nondeterministic Turing machine $M$ and $\overline{L}(M)$ accepted by ${\rm co}\mathcal{NP}$ machine $M$, we will apply only simulation techniques. So, we first try to show that all polynomial-time nondeterministic Turing machines are enumerable. Then, we use a universal nondeterministic Turing machine to simulate all polynomial-time nondeterministic Turing machines, creating a language that is not in ${\rm co}\mathcal{NP}$. To stress it is “novel,” it is because this method is the first appearing in this paper, and we have never seen such an attempt appeared in previous works or such an idea in the relevant standard textbooks, such as [AHU74, AB09, Sip13, Pap94].
@@ 1.4. Organization
The remainder of this paper is organized as follows: For the convenience of the reader, we will review some notions in the next section closely associated with our discussions and fix some notation we will use in the following context, with some useful technical lemmas also presented in the next section. In Section 3, we prove that all polynomial-time nondeterministic Turing machines are enumerable and all ${\rm co}\mathcal{NP}$-machines are enumerable, thus showing that there exists an enumeration of all ${\rm co}\mathcal{NP}$-machines. Section 4 contains the definition of our universal nondeterministic Turing machine, which accepts a language $L_{s}$ not in ${\rm co}\mathcal{NP}$, showing the desired lower bounds. And the Section 5 is devoted to showing that the language $L_{s}$ is in fact in $\mathcal{NP}$, proving the upper bounds required, in which we finish the proof of Theorem 1.1 at our own convenience. In Section 6 we will show Theorem 1.2 to break the so-called Relativization Barrier. In Section 7, we show that there are ${\rm co}\mathcal{NP}$-intermediate languages in the complexity class ${\rm co}\mathcal{NP}$. In Section 8, we prove the Theorem 1.8 which says that no Frege proof systems are polynomially bounded, thus answering an important open question in the area of proof complexity. Finally, we list some open problems for future research in the last section.
@@ 2. Preliminaries
Let $\mathbb{N}_{1}=\{1,2,3,\cdots\}$ where $+\infty\notin\mathbb{N}_{1}$, i.e., the set of all positive integers. We also denote by $\mathbb{N}$ the set of natural numbers, i.e.,
@@
Let $\Sigma$ be an alphabet, for finite words $w,v\in\Sigma^{*}$, the concatenation of $w$ and $v$, denoted by $w\circ v$, is $wv$. For example, suppose that $\Sigma=\{0,1\}$, and $w=100001$, $v=0111110$, then
@@
The little $o$ notation also indicates the order of growth of some quantity as a function of $n$ or the limiting behavior of a function but with different meaning. Specifically, that $T(n)$ is little $o$ of $t(n)$, i.e.,
@@
A nondeterministic Turing machine $M$ works in time $t(n)$ (or of time complexity $t(n)$), if for any input $w\in I^{*}$ where $I$ is the input alphabet of $M$, $M(w)$ will halt within $t(|w|)$ steps.
@@
By default, a word $w$ is accepted by a polynomial-time (say, $t(n)$ time-bounded) nondeterministic Turing machine $M$ if there exists at least one computation path $\gamma$ putting the machine into an accepting state (i.e., stopping in the “accepting” state) on input $w$ (in such a situation, we call $M$ accepts $w$). Specifically, $M$ accepts the language $L(M)$, then $w\in L(M)$ if and only if $M$ accepts $w$. In other words, there exists at least one accepting path for $M$ on input $w$. This is the “there exists at least one accepting path” condition for nondeterministic Turing machines, based on which the complexity class ${\rm NTIME}[t(n)]$ is defined:
@@
A ${\rm co}\mathcal{NP}$ machine $M$ is exactly a polynomial-time nondeterministic Turing machine $M$ working under the “there exists no accepting path” condition. We define the language accepted by ${\rm co}\mathcal{NP}$ machine $M$ to be $\overline{L}(M)$.222 In short, there is an interesting relationship: ${\rm co}\mathcal{NP}$ machine $M$ accepts the language $\overline{L}(M)$ but polynomial-time nondeterministic Turing machine $M$ rejects $\overline{L}(M)$; polynomial-time nondeterministic Turing machine $M$ accepts the language $L(M)$ but ${\rm co}\mathcal{NP}$ machine $M$ rejects $L(M)$. Namely, for any input $w$, $w\in\overline{L}(M)$ if and only if polynomial-time nondeterministic Turing machine $M$ rejects $w$.
@@
In fact, we can define the complexity class ${\rm co}\mathcal{NP}$ in a equivalent way using polynomial-time deterministic Turing machines as verifiers and “for all” witnesses, given as follows:
@@ 3. All Polynomial-time Nondeterministic Turing Machines Are Enumerable
Following Definition 2.2, a polynomial-time nondeterministic Turing machine, say a $n^{k}+k$ time-bounded nondeterministic Turing machine, can be represented by the tuple of $(M,k)$, where $M$ is the $n^{k}+k$ time-bounded nondeterministic Turing machine itself and $k$ is the degree of polynomial $|x|^{k}+k$ such that $M(x)$ will halt within $|x|^{k}+k$ steps for any input $x$ of length $|x|$. We call such a positive integer $k$ the order of $(M,k)$.
@@
Therefore, we have reached the following important result:
@@
There exists a one-one correspondence between $\mathbb{N}_{1}$ and the set $\{(M,k)\}$ of all polynomial-time nondeterministic Turing machines. In other words, the set $\{(M,k)\}$ of all polynomial-time nondeterministic Turing machines is enumerable (i.e., all polynomial-time nondeterministic Turing machines are in the list of $e$), and all languages in $\mathcal{NP}$ are enumerable (with languages appearing multiple times).
@@
$\pi:\mathbb{N}\times\mathbb{N}\rightarrow\mathbb{N}$
@@
4. Simulation of All Polynomial-time Nondeterministic Turing Machines
In computational complexity, the most basic approach to show hierarchy theorems uses simulation and diagonalization, because the simulation and diagonalization technique [Tur37, HS65] is a standard method to prove lower bounds on uniform computing models (i.e., the Turing machine model); see for example [AHU74]. These techniques work well in deterministic time and space measures; see e.g. [For00, FS07]. However, they do not work for nondeterministic time, which is not known to be closed under complement at present — one of the main issues discussed in this paper; hence it is unclear how to define a nondeterministic machine that “does the opposite”; see e.g. [FS07].
For nondeterministic time, we can still do a simulation but can no longer negate the answer directly. In this case, we apply the lazy diagonalization to show the nondeterministic time hierarchy theorem; see e.g. [AB09]. It is worth noting that Fortnow [For11] developed a much more elegant and simple style of diagonalization to show nondeterministic time hierarchy; see also [FS10]. Generally, lazy diagonalization [Žák83, AB09, FS07] is a clever application of the standard diagonalization technique [Tur37, HS65, AHU74]. The basic strategy of lazy diagonalization is that for any $n\leq j<m=n^{t(n)}$, $M$ on input $1^{j}$ simulates $M_{i}$ on input $1^{j+1}$, accepting if $M_{i}$ accepts and rejecting if $M_{i}$ rejects. When $j=m$, $M$ on input $1^{j}$ simulates $M_{i}$ on input $x$ deterministically, accepting if $M_{i}$ rejects and rejecting if $M_{i}$ accepts. Since $m=n^{t(n)}$, $M$ has enough time to perform the trivial exponential-time deterministic simulation of $M_{i}$ on input $1^{n}$; what we are doing is deferring the diagonalization step by linking $M$ and $M_{i}$ together on larger and larger inputs until $M$ has an input large enough that it can actually “do the opposite” deterministically; see e.g. [FS07] for details.
However, although all ${\rm co}\mathcal{NP}$ machines are proved to be enumerable and the enumeration is $e$ given in Section 3, it is hard to apply the standard diagonalization technique or the lazy-diagonalization technique from [Žák83, AB09, FS07] to create a language that is not in ${\rm co}\mathcal{NP}$. Instead, our main goal of this section is to apply simulation techniques rather than diagonalization techniques and the particular relations between $L(M)$ accepted by a polynomial-time nondeterministic Turing machine $M$ and $\overline{L}(M)$ accepted by a ${\rm co}\mathcal{NP}$ machine $M$ to show that there is a language $L_{s}$ accepted by a universal nondeterministic Turing machine with the property that $L_{s}\notin{\rm co}\mathcal{NP}$.
@@ 4. Simulation of All Polynomial-time Nondeterministic Turing Machines
Let $U$ be a four-tape nondeterministic Turing machine which operates as follows on an input string $x$ of length of $n$.
@@
    $U$ decodes the tuple encoded by $x$. If $x$ is not the encoding of some ${\rm co}\mathcal{NP}$ machine $M_{l}$ for some $l$ (by Definition 2.4, equivalently, $x$ can not be decoded into some polynomial-time nondeterministic Turing machine, i.e., $x$ represents the trivial Turing machine with an empty next-move function) then GOTO $5$, else determines $t$, the number of tape symbols used by $M_{l}$; $s$, its number of states; and $k$, its order. The third tape of $U$ can be used as “scratch” memory to calculate $t$.
@@
    cells each, the blocks being separated by single cell holding a marker ##\##, i.e., there are
@@
    time and output its answer, i.e., accepting if $M_{l}$ accepts and rejecting if $M_{l}$ rejects.”
@@
    Since $x$ is encoding of the trivial Turing machine with an empty next-move function. Then $U$ sets up a block of
@@
    steps, so is within $O(n)$. Note that the number of $2$ in
@@
(4.1)4.1( 4.1 ) $1^{|w|-m}\langle M_{i}\rangle\in L_{s}\quad\text{if and only if}\quad 1^{|w|-m}\langle M_{i}\rangle\in L(M_{w})=L(M_{i}).$
@@
From (4.1) and (4.2) we conclude that if $U$ accepts the input $1^{|w|-m}\langle M_{i}\rangle$, i.e.,
$1^{|w|-m}\langle M_{i}\rangle\in L_{s}$
@@
$1^{|w|-m}\langle M_{i}\rangle\notin L_{s}$
@@
One may wish to apply simulation techniques to the case of $\mathcal{P}$ versus ${\rm co}\mathcal{P}$ and doubt that simulation techniques are unable to produce a language not in ${\rm co}\mathcal{P}$. In this scenario, we remark that the simulation technique is unable to create a language not in ${\rm co}\mathcal{P}$ via a universal deterministic Turing machine $U^{\prime}$, because a polynomial-time deterministic Turing machine ($\mathcal{P}$ machine) that works under the “there exists no accepting path” condition is still a polynomial-time deterministic Turing machine. Namely, the “there exists at least one accepting path” condition and the “there exists no accepting path” condition for deterministic Turing machines are symmetrical. Specifically, for an arbitrary deterministic Turing machine $M$, if $M$ accepts the input $w$, then there is only one accepting path, and if $M$ rejects the input $w$, then there is also only one rejecting path. This is different from nondeterministic Turing machines. In short, if we apply the same technique, i.e., the simulation technique, to create a language not in ${\rm co}\mathcal{P}$, then the universal deterministic Turing machine $U^{\prime}$ (it runs within time $O(n^{k})$ for any $k\in\mathbb{N}_{1}$) needed to be constructed accepts the language $L(U^{\prime})\in\mathcal{P}$ (the proof is similar to the case in Section 5 below). But by the symmetrical properties between the “there exists at least one accepting path” condition and the “there exists no accepting path” condition with respect to deterministic Turing machines, we clearly have that $L(U^{\prime})\in{\rm co}\mathcal{P}$. Thus, we are unable to claim that $L(U^{\prime})\notin{\rm co}\mathcal{P}$.

### v19

@@ 1.3. Our Approach
However, as a novel idea and attempt, we will apply simulation techniques to this problem. To the best of our knowledge, if we want to create a language not in a specific complexity class $\mathcal{C}$, the most obvious way is to apply the diagonalization technique or the lazy-diagonalization technique, because the simulation and diagonalization technique [Tur37, HS65] is a standard method (currently the only way) to prove lower bounds on uniform computing models (i.e., the Turing machine model). But, we find that it is hard to apply standard diagonalization or lazy-diagonalization techniques to produce a language that is not in ${\rm co}\mathcal{NP}$. Instead, by the relation between the language $L(M)$ accepted by polynomial-time nondeterministic Turing machine $M$ and $\overline{L}(M)$ accepted by ${\rm co}\mathcal{NP}$ machine $M$, we will apply only simulation techniques. So, we first try to show that all polynomial-time nondeterministic Turing machines are enumerable. Then, we use a universal nondeterministic Turing machine to simulate all polynomial-time nondeterministic Turing machines, creating a language that is not in ${\rm co}\mathcal{NP}$. To stress that it is “novel,” it is because this method is the first appearing in this paper, and we have never seen such an attempt appear in previous works or such an idea in the relevant standard textbooks, such as [AHU74, AB09, Sip13, Pap94].
@@ 1.4. Organization
The remainder of this paper is organized as follows: For the convenience of the reader, we will review some notions in the next section closely associated with our discussions and fix some notation we will use in the following context, with some useful technical lemmas also presented in the next section. In Section 3, we prove that all polynomial-time nondeterministic Turing machines are enumerable and all ${\rm co}\mathcal{NP}$-machines are enumerable, thus showing that there exists an enumeration of all ${\rm co}\mathcal{NP}$-machines. Section 4 contains the definition of our universal nondeterministic Turing machine, which accepts a language $L_{s}$ not in ${\rm co}\mathcal{NP}$, showing the desired lower bounds. The Section 5 is devoted to showing that the language $L_{s}$ is in fact in $\mathcal{NP}$, proving the upper bounds required, in which we finish the proof of Theorem 1.1 at our own convenience. In Section 6 we will show Theorem 1.2 to break the so-called Relativization Barrier. In Section 7, we show that there are ${\rm co}\mathcal{NP}$-intermediate languages in the complexity class ${\rm co}\mathcal{NP}$. In Section 8, we prove Theorem 1.8, which says that no Frege proof systems are polynomially bounded, thus answering an important open question in the area of proof complexity. Finally, we list some open problems for future research in the last section.
@@ 2. Preliminaries
Let $\mathbb{N}_{1}=\{1,2,3,\cdots\}$, where $+\infty\notin\mathbb{N}_{1}$, i.e., the set of all positive integers. We also denote by $\mathbb{N}$ the set of natural numbers, i.e.,
@@
Let $\Sigma$ be an alphabet. For finite words $w,v\in\Sigma^{*}$, the concatenation of $w$ and $v$, denoted by $w\circ v$, is $wv$. For example, suppose that $\Sigma=\{0,1\}$, and $w=100001$, $v=0111110$; then
@@
The little $o$ notation also indicates the order of growth of some quantity as a function of $n$ or the limiting behavior of a function, but with different meaning. Specifically, that $T(n)$ is little $o$ of $t(n)$, i.e.,
@@
A nondeterministic Turing machine $M$ works in time $t(n)$ (or of time complexity $t(n)$) if, for any input $w\in I^{*}$, where $I$ is the input alphabet of $M$, $M(w)$ will halt within $t(|w|)$ steps.
@@
By default, a word $w$ is accepted by a polynomial-time (say, $t(n)$ time-bounded) nondeterministic Turing machine $M$ if there exists at least one computation path $\gamma$ putting the machine into an accepting state (i.e., stopping in the “accepting” state) on input $w$ (in such a situation, we say $M$ accepts $w$). Specifically, if $M$ accepts the language $L(M)$, then $w\in L(M)$ if and only if $M$ accepts $w$. In other words, there exists at least one accepting path for $M$ on input $w$. This is the “there exists at least one accepting path” condition for nondeterministic Turing machines, based on which the complexity class ${\rm NTIME}[t(n)]$ is defined:
@@
A ${\rm co}\mathcal{NP}$ machine $M$ is exactly a polynomial-time nondeterministic Turing machine $M$ working under the “there exists no accepting path” condition. We define the language accepted by the ${\rm co}\mathcal{NP}$ machine $M$ to be $\overline{L}(M)$. 222 In short, there is an interesting relationship: ${\rm co}\mathcal{NP}$ machine $M$ accepts the language $\overline{L}(M)$, but polynomial-time nondeterministic Turing machine $M$ rejects $\overline{L}(M)$; polynomial-time nondeterministic Turing machine $M$ accepts the language $L(M)$, but ${\rm co}\mathcal{NP}$ machine $M$ rejects $L(M)$. Namely, for any input $w$, $w\in\overline{L}(M)$ if and only if polynomial-time nondeterministic Turing machine $M$ rejects $w$.
@@
In fact, we can define the complexity class ${\rm co}\mathcal{NP}$ in an equivalent way using polynomial-time deterministic Turing machines as verifiers and “for all” witnesses, given as follows:
@@ 3. All Polynomial-time Nondeterministic Turing Machines Are Enumerable
Following Definition 2.2, a polynomial-time nondeterministic Turing machine, say an $n^{k}+k$ time-bounded nondeterministic Turing machine, can be represented by the tuple $(M,k)$, where $M$ is the $n^{k}+k$ time-bounded nondeterministic Turing machine itself and $k$ is the degree of polynomial $|x|^{k}+k$ such that $M(x)$ will halt within $|x|^{k}+k$ steps for any input $x$ of length $|x|$. We call such a positive integer $k$ the order of $(M,k)$.
@@
Therefore, we have arrived at the following important result:
@@
There exists a one-to-one correspondence between $\mathbb{N}_{1}$ and the set $\{(M,k)\}$ of all polynomial-time nondeterministic Turing machines. In other words, the set $\{(M,k)\}$ of all polynomial-time nondeterministic Turing machines is enumerable (i.e., all polynomial-time nondeterministic Turing machines are in the list of $e$), and all languages in $\mathcal{NP}$ are enumerable (with languages appearing multiple times).
@@
$\pi:\mathbb{N}\times\mathbb{N}\rightarrow\mathbb{N},$
@@
4. Simulating of All Polynomial-time Nondeterministic Turing Machines
In computational complexity, the most basic approach to show hierarchy theorems uses simulation and diagonalization, because the simulation and diagonalization technique [Tur37, HS65] is a standard method to prove lower bounds on uniform computing models (i.e., the Turing machine model); see for example [AHU74]. These techniques work well in deterministic time and space measures; see e.g., [For00, FS07]. However, they do not work for nondeterministic time, which is not known to be closed under complement at present, i.e., one of the main issues discussed in this paper; hence it is unclear how to define a nondeterministic machine that “does the opposite”; see e.g., [FS07].
For nondeterministic time, we can still do a simulation but can no longer negate the answer directly. In this case, we apply the lazy diagonalization to show the nondeterministic time hierarchy theorem; see e.g. [AB09]. It is worth noting that Fortnow [For11] developed a much more elegant and simple style of diagonalization to show nondeterministic time hierarchy; see also [FS10]. Generally, lazy diagonalization [Zak83, AB09, SFM78] is a clever application of the standard diagonalization technique [Tur37, HS65, AHU74]. The basic strategy of lazy diagonalization is that for any $n\leq j<m=n^{t(n)}$, $M$ on input $1^{j}$ simulates $M_{i}$ on input $1^{j+1}$, accepting if $M_{i}$ accepts and rejecting if $M_{i}$ rejects. When $j=m$, $M$ on input $1^{j}$ simulates $M_{i}$ on input $x$ deterministically, accepting if $M_{i}$ rejects and rejecting if $M_{i}$ accepts. Since $m=n^{t(n)}$, $M$ has enough time to perform the trivial exponential-time deterministic simulation of $M_{i}$ on input $1^{n}$; what we are doing is deferring the diagonalization step by linking $M$ and $M_{i}$ together on larger and larger inputs until $M$ has an input large enough that it can actually “do the opposite” deterministically; see e.g. [FS07] for details.
However, although the set of all ${\rm co}\mathcal{NP}$ machines is proved to be enumerable and the enumeration is $e$ given in Section 3, it is challenging (even impossible) to apply the standard diagonalization technique or the lazy-diagonalization technique from [Zak83, AB09, FS07, SFM78] to create a language that is not in ${\rm co}\mathcal{NP}$. Instead, our main goal of this section is to apply simulation techniques rather than diagonalization techniques and the particular relations between $L(M)$ accepted by a polynomial-time nondeterministic Turing machine $M$ and $\overline{L}(M)$ accepted by a ${\rm co}\mathcal{NP}$ machine $M$ to show that there is a language $L_{s}$ accepted by a universal nondeterministic Turing machine with the property that $L_{s}\notin{\rm co}\mathcal{NP}$.
@@ 4. Simulating of All Polynomial-time Nondeterministic Turing Machines
Let $U$ be a four-tape nondeterministic Turing machine that operates as follows on an input string $x$ of length $n$.
@@
    $U$ decodes the tuple encoded by $x$. If $x$ is not the encoding of some ${\rm co}\mathcal{NP}$ machine $M_{l}$ for some $l$ (by Definition 2.4, equivalently, $x$ can not be decoded into some polynomial-time nondeterministic Turing machine, i.e., $x$ represents the trivial Turing machine with an empty next-move function), then GOTO $5$, else determine $t$, the number of tape symbols used by $M_{l}$; $s$, its number of states; and $k$, its order. The third tape of $U$ can be used as “scratch” memory to calculate $t$.
@@
    cells each, the blocks being separated by a single cell holding a marker ##\##, i.e., there are
@@
    time and outputs its answer, i.e., accepting if $M_{l}$ accepts and rejecting if $M_{l}$ rejects.”
@@
    Since $x$ is an encoding of the trivial Turing machine with an empty next-move function. Then $U$ sets up a block of
@@
    steps, so it is within $O(n)$. Note that the number of $2$ in
@@
(4.1)4.1( 4.1 ) $1^{|w|-m}\langle M_{i}\rangle\in L_{s}\quad\text{if and and only if}\quad 1^{|w|-m}\langle M_{i}\rangle\in L(M_{w})=L(M_{i}).$
@@
From (4.1) and (4.2), we conclude that if $U$ accepts the input $1^{|w|-m}\langle M_{i}\rangle$, i.e.,
$1^{|w|-m}\langle M_{i}\rangle\in L_{s},$
@@
$1^{|w|-m}\langle M_{i}\rangle\notin L_{s},$
@@
One may wish to apply simulation techniques to the case of $\mathcal{P}$ versus ${\rm co}\mathcal{P}$ and doubt that simulation techniques are unable to produce a language not in ${\rm co}\mathcal{P}$. In this scenario, we remark that the simulation technique is unable to create a language not in ${\rm co}\mathcal{P}$ via a universal deterministic Turing machine $U^{\prime}$, because a polynomial-time deterministic Turing machine (we call it a $\mathcal{P}$ machine) that works under the “there exists no accepting path” condition is still a polynomial-time deterministic Turing machine. Namely, the “here exists at least one accepting path” condition and the “there exists no accepting path” condition for deterministic Turing machines are symmetrical. Specifically, for an arbitrary deterministic Turing machine $M$, if $M$ accepts the input $w$, then there is only one accepting path, and if $M$ rejects the input $w$, then there is also only one rejecting path. This is different from nondeterministic Turing machines. In short, if we apply the same technique, i.e., the simulation technique, to create a language not in ${\rm co}\mathcal{P}$, then the universal deterministic Turing machine $U^{\prime}$ (it runs within time $O(n^{k})$ for any $k\in\mathbb{N}_{1}$) needed to be constructed accepts the language $L(U^{\prime})\in\mathcal{P}$ (the proof is similar to the case in Section 5 below). But by the symmetrical properties between the “there exists at least one accepting path” condition and the “there exists no accepting path” condition with respect to deterministic Turing machines, we clearly have that $L(U^{\prime})\in{\rm co}\mathcal{P}$. Thus, we are unable to claim that $L(U^{\prime})\notin{\rm co}\mathcal{P}$.